OpenStudy (anonymous):

HELP! WILL MEDAL AND FAN!In a survey, four out of seven people named blue as their favorite color. One out of three named red. If 1092 people were included in the survey, how many named neither red or blue. I need help how to solve this problem. I will medal and fan.

OpenStudy (anonymous):

about 70%

Directrix (directrix):

4/7 of those surveyed chose blue: 4/7 of 1092 = 624 blues 1/3 of those surveyed chose red: 1/3 of 1092 = 364 reds Reds + Blues = 624 + 364 = ? @tigerlilly_1 Subtract that number from 1092 to get the number of people who chose neither blue nor red. Post what you get, okay?

OpenStudy (anonymous):

104

Directrix (directrix):

That is what I got, too.

OpenStudy (anonymous):

Thank you so much!